An employee at See's Candy Store needs a special mixture of candy. She has creme-filled chocolates that sell for $3.00 per pound, chocolate-covered nuts that sell for $2.70 per pound, and chocolate-covered raisins that sell for $2.25 per pound. She wants to have twice as many raisins as nuts in the mixture. Find how many pounds of each she should use to make 45 pounds worth $2.80 per pound.
--------------------------------------------------

Here's how I set up the problem:
a + b + c = 45
3a + 2.7b + 2.25c = 2.8(45)
Eventually, I ended up with no solution, which I am pretty certain isn't the correct answer. My textbook does not have any example problems of this nature, so I don't really have any idea as to how to go about solving them; any help would be greatly appreciated.

Answer by richwmiller(17219)

you have three unknowns
you need at least three equations
c for creme
n for nuts
r for raisins
c+n+r=45
3c + 2.7n + 2.25r= 2.8(45)
Now you need a third equation.
r=2n
c=30., n=5., r=10.